Powder milk is packed in 1-kilogram bag. An inspector from the Department of Trade and Industry (DTI) suspects that bags may not contain 1 kilogram. A sample of 40 bags produces a mean of 0.96 kilograms and a standard deviation of 0.12 kilogram. Is there enough evidence to conclude that the bags do not contain 1 kilogram as stated at 0.05 level of confidence?


1
Expert's answer
2021-09-29T07:29:25-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\geq1"

"H_1:\mu<1"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05,"

"df=n-1=40-1=39" degrees of freedom, and and the critical value for a left-tailed test is "t_c= -1.684875."

The rejection region for this left-tailed test is "R = \{t: t < -1.684875\}."

The t-statistic is computed as follows:


"t-=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{0.96-1}{0.12/\sqrt{40}}=-2.108185"

Since it is observed that "t = -2.108185 < -1.684875=t_c ," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value for left-tailed "\alpha=0.05, df=39," "t=-2.108185" is "p=0.020746," and since "p=0.020746<0.05=\alpha,"

it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"  is less than "1," at the "\alpha = 0.05" significance level.

Therefore, there is enough evidence to claim that the bags do not contain 1 kilogram as stated, at the "\alpha = 0.05" significance level.